I do not understand this at all. I know that I have to substitute the x and y values but I don't know how to solve it. It is solving systems using substitution. One of the problems is: y=3x-10 and the other equation that goes with it is y=2x-5
Answer by Greenfinch(383)

These are simultaneous equations, which means that they can be solved if they are both true AT THE SAME TIME. This can either be done by getting rid of the x or y, or, in this case by substitution.
If y = 3x - 10 and y = 2x - 5. then
3x - 10 = 2x - 5 because they both equal y and we can substitute 2x -5 for y
This leads to x = 5 ( 3x - 2x and -5 -(-10))
Put that in one of the original equations and we get 3x5 -10 = 5 or 2 -5 = 5
So the answers are x=5 and y=5